Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-11-24

Observers: Z.Kostrzewa-Rutkowska, K.Lopez
TAT: J. Chen, M. Fraser, J. Sollerman
Support team: P. Jonker
Instruments: EFOSC2

Weather conditions: 
evening UT 0:30: thin clouds, humidity 48%
morning UT 8:15: clear, humidity 16%

Calibrations:
EFOSC2 bias, dome flats, arcs daytime
EFOSC2 evening skyflats BVr, morning BVri 

LOG:

UT      Target           Type       Filter/Grism ExpTime[s]  Slit   Seeing  Airmass    Comments
00:30                                                                                  System issues
00:46   Feige110         STD        gr11/13/16              1.0/1.5 2.0      1.15
01:05   AT2017gpp        Followup   gr13            2700    1.5     2.0      1.20      Aborted - target not visible
01:15   ATLAS17nja       Class      gr13            1500    1.0     1.6      1.09
01:50   AT2017hxz        Followup   gr13            2700    1.5     2.2      1.19
02:45   ATLAS17njg       Class      gr13            1500    1.5     2.1      1.25      Problem with focus at the beginning
03:20   L745a            STD        gr11/13/16              1.0/1.5 1.8      2.07
03:46   AT2017hbj        Followup   gr13            2700    1.5     2.1      1.23   
04:38   PS17fcj          Class      gr13            900     1.5     1.6      1.35
05:08   MASTERJ034412..  Class?     gr11            600     1.5     2.1      1.34
05:25   PS17fcg          Class      gr13            1500    1.5     1.9      1.66
06:00   ATLAS17ngb       Class      gr13            600     1.5     1.8      1.30
06:20   AT2017ckq        Followup   gr13            2700    1.0     1.1      1.95
07:15   AT2017ckq        Followup   gr13            2700    1.0     1.0      1.52
08:10   L745a            STD        gr11/13/16              1.0/1.5 1.3      1.02

   
     

Comments